# Embodied Navigation Foundation Model

### v1

@@ Abstract
Navigation is a fundamental capability in embodied AI, representing the intelligence required to perceive and interact within physical environments following language instructions. Despite significant progress in large Vision-Language Models (VLMs), which exhibit remarkable zero-shot performance on general vision-language tasks, their generalization ability in embodied navigation remains largely confined to narrow task settings and embodiment-specific architectures. In this work, we introduce a cross-embodiment and cross-task Navigation Foundation Model (NavFoM), trained on eight million navigation samples that encompass quadrupeds, drones, wheeled robots, and vehicles, and spanning diverse tasks such as vision-and-language navigation, object searching, target tracking, and autonomous driving. NavFoM employes a unified architecture that processes multimodal navigation inputs from varying camera configurations and navigation horizons. To accommodate diverse camera setups and temporal horizons, NavFoM incorporates identifier tokens that embed camera view information of embodiments and temporal context of tasks. Furthermore, to meet the demands of real-world deployment, NavFoM controls all observation tokens using a dynamically adjusted sampling strategy under a limited token length budget. Extensive evaluations on public benchmarks demonstrate that our model achieves state-of-the-art or highly competitive performance across multiple navigation tasks and embodiments without requiring task-specific fine-tuning. Real-world experiments further confirm the strong generalization capability and practical applicability of our approach.
@@ 1 Introduction
In this work, we toward building a cross-task and cross-embodiment embodied navigation foundation model (Figure 1), NavFoM, trained on eight million navigation samples spanning diverse embodiments and tasks. Inspired by humans’ ability to accomplish a wide range of navigation tasks primarily through visual sensory input and the recent success of vision-only navigation methods (Shah et al., 2023a; Zeng et al., ), we formulate the generalist navigation task as processing egocentric videos (captured by one or more cameras mounted on the robot) alongside language instructions, and predicting subsequent trajectories to fulfill those instructions. This formulation is compatible with most existing navigation task settings (Contributors, 2023; Wang et al., 2024a).
To align generalizable embodiments across diverse camera configurations, we introduce temporal-viewpoint indicator tokens (TVI tokens) to identify both the viewpoint of camera setups and the temporal information of the navigation horizon. By dynamically adjusting these TVI tokens, our method enables co-tuning across different camera setups and supports joint training with both image-QA and video-QA samples (Shen et al., 2024; Li et al., 2023). Furthermore, to address the demands of practical deployment, we propose a token Budget-Aware Temporal Sampling (BATS) strategy, which dynamically samples navigation history tokens based on a forgetting curve constrained by a token budget. This token sampling approach balances performance and inference speed, enhancing the practicality of our method for real-world deployment.
@@
Our experiments demonstrate that NavFoM achieves substantial advancements in generalist navigation. Without task-specific fine-tuning, NavFoM attains state-of-the-art or competitive performance (Figure 2) across diverse public benchmarks for a variety of embodiments. On VLN-CE RxR (Ku et al., 2020), NavFoM improves performance in multi-camera settings (from $56.3\%$ to $64.4\%$ SR) and in single-camera settings (from $51.8\%$ to $57.4\%$ SR) compared to prior baselines. On HM3D-OVON (Yokoyama et al., 2024b), our method achieves $45.2\%$ SR in a zero-shot setting, outperforming the previous fine-tuned SOTA method ($43.6\%$ SR). Similarly strong results are observed across various benchmarks in object searching, tracking, and autonomous driving. We further validate NavFoM through real-world experiments on multiple robotic platforms, including humanoid robots, quadrupeds, drones, and wheeled robots. These results underscore its strong generalizability and highlight promising progress toward generalist navigation.
@@ 2 Related Works
Large models for Navigation. Integrating Large Models (LLMs and VLMs) into robotic navigation has shifted the field from traditional learning-based methods toward leveraging pre-trained knowledge for open-world understanding and strong generalization capability (Deng et al., 2025; Zhou et al., 2025a; 2024a). One straightforward approach (Zhou et al., 2023; Shah et al., 2022; Qiao et al., 2023) is to use off-the-shelf LLMs in a zero-shot manner. These works emphasize interpretability via chain-of-thought mechanisms (Pan et al., 2023; Long et al., 2023; Lin et al., 2025) and structured reasoning frameworks (Chen et al., 2024b; Qiao et al., 2025). However, abstracting dense visual information into text results in sparse environmental observations and is limited to static environments. Another pathway (Cheng et al., 2025; Zhang et al., 2024a; 2025a; Wei et al., 2025; Wang et al., 2025c) involves end-to-end fine-tuning of video-based or image-based (Zhou et al., 2025b; Zheng et al., 2024; Zhang et al., 2025c) vision-language models with navigation data to enable VLMs to master navigation capabilities. However, existing methods mostly focus on isomorphic embodiments, overlooking the potential training synergy between different embodiments and tasks. In this work, we make an initial effort to extend the navigation policy to a broader domain of cross-embodiment and cross-task navigation.
@@
Cross-embodiments Navigation. The development of navigation models that generalize across diverse embodiments—varying in shape, size, and sensor configurations—remains a significant challenge in embodied AI. Recent efforts (Shah et al., 2023a; b; Yang et al., 2024b; Wang et al., 2020; Eftekhar et al., 2024; Hirose et al., 2023; Putta et al., 2024; Curtis et al., 2024; Wang et al., 2025a; Zhang et al., 2025b; Geng et al., 2025) have demonstrated the potential of transformer-based policies trained on large-scale, cross-embodiment datasets to achieve robust performance across various robotic platforms without the need for manual alignment of observation and action spaces. However, these models often process multimodal inputs without incorporating explicit spatial and temporal cues, which can lead to ambiguities due to the differing geometric interpretations of data from various embodiments. This approach may also result in data inefficiencies and limited generalization to out-of-distribution embodiments. (Eftekhar et al. (2024); Wang et al. (2025b)) In contrast, NavFoM introduces temporal-spatial indicator tokens that encode the observation configurations, enabling the model to better interpret multimodal inputs across different embodiments.
@@
Cross-task Navigation. Recent advances in embodied AI (O’Neill et al., 2024; Team et al., 2024; Kim et al., ; Bjorck et al., 2025; Black et al., 2024; Intelligence et al., 2025; Bu et al., 2025b; a; Qu et al., 2025) demonstrate that generalist models built upon foundation models can effectively transfer knowledge across diverse tasks. In the domain of navigation, prior studies (Zhou et al., 2024b; Wang et al., 2022; Long et al., 2024; Song et al., 2025; Zhang et al., 2025a; Gao et al., 2025; Yin et al., 2025; Ruan et al., 2025) have shown that integrating data from different categories of navigation tasks can lead to stronger performance across various navigation scenarios. An early work, VIENNA (Wang et al., 2022), leveraged reinforcement learning to train an agent in simulators. More recently, Uni-Navid (Zhang et al., 2025a) developed a generalist model based on video-based vision-language models (VLMs) that learns generalized navigation skills through cross-task learning across four types of tasks: vision-and-language navigation, object-goal navigation, embodied question answering, and human following. However, these approaches are limited to constrained settings (e.g., indoor controllable environments), whereas our work extends to broader scenarios (e.g., including autonomous driving and UAV navigation) and unifies all tasks under a general formulation. In this framework, the model takes RGB videos and natural language instructions as input and outputs an executable trajectory.
@@ 3.2 Data
Vision-and-Language Navigation (3.37 M) requires an agent to interpret natural language instructions and egocentric visual observations, align the instructions with visual inputs, and plan subsequent actions to reach described landmarks. Following a broad definition of VLN (Zheng et al., 2024; Wang et al., 2025b; Zhou et al., 2024b), we consider both indoor environments (e.g., VLN-CE on R2R (Krantz et al., 2020) and RxR (Ku et al., 2020)) and outdoor environments (e.g., OpenUAV (Wang et al., 2024a)), deployed on robots and unmanned aerial vehicles (UAVs), respectively. For VLN-CE on R2R and RxR (2.94 M), we capture multi-view RGB videos, annotated instructions, and trajectory data while the robot follows the ground-truth path. The multi-view RGB setup consists of a fixed front-view camera and randomly sampled surrounding cameras (ranging from one to eight). Camera heights are randomized between 0.6 m and 1.5 m, and the horizontal fields of view (HFoV) vary between 75 and 120. For the OpenUAV dataset (429 K), we record camera streams from the front, left, right, and rear views for all sequences. Other randomization strategies remain consistent with those used in the VLN-CE tasks.
@@ 4.1 Experiment Setup
Metrics. To evaluate navigation performance, we follow the standard evaluation metrics (Anderson et al., 2018; Zhang et al., 2023), including success rate (SR), oracle success rate (OS), success weighted by path length (SPL), normalized Dynamic Time Warping (nDTW) and navigation error from goal (NE). Espeecially for tracking task, we leverage tracking rate (TR) (Puig et al., 2023), which measure the ratio of success tracking steps over all steps. Note that the success criteria change among different navigation tasks, we therefore use the default success criteria of each benchmark. For autonomous driving evaluation, we report L2 distance and Collision Rate (CR) for open-loop planning (Caesar et al., 2020a). For closed-loop evaluation on NAVSIM, we use the PDM score (PDMS) (Dauner et al., 2024b), which is a weighted combination of several sub-metrics: no at-fault collisions (NC), drivable area compliance (DAC), time-to-collision (TTC), comfort (Comf.), and ego progress (EP).

### v2

@@ Abstract
Navigation is a fundamental capability in embodied AI, representing the intelligence required to perceive and interact within physical environments. To achieve such intelligence, recent advanced works leverage Vision-Language Models (VLMs), which demonstrate strong generalizability and possess a well-suited formulation for navigation. However, these approaches remain largely confined to narrow task settings and embodiment-specific architectures. In this work, we introduce a cross-embodiment and cross-task Navigation Foundation Model (NavFoM), trained on eight million navigation samples that encompass quadrupeds, drones, wheeled robots, and vehicles, and spanning diverse tasks such as vision-and-language navigation, object searching, target tracking, and autonomous driving. NavFoM employs a unified architecture that processes multimodal navigation inputs from varying camera configurations and navigation horizons. To accommodate diverse camera setups and temporal horizons, NavFoM incorporates identifier tokens that embed camera view information of embodiments and the temporal context of tasks. Furthermore, to meet the demands of real-world deployment, NavFoM controls all observation tokens using a dynamically adjusted sampling strategy under a limited token length budget. Extensive evaluations on seven public benchmarks demonstrate that our model achieves state-of-the-art or highly competitive performance across different navigation tasks and embodiments without requiring task-specific fine-tuning. Additional real-world experiments further confirm the strong generalizability and practical applicability of our approach.
@@ 1 Introduction
In this work, we toward building a cross-task and cross-embodiment embodied navigation foundation model, NavFoM, trained on eight million navigation samples spanning diverse embodiments and tasks. Inspired by humans’ ability to accomplish a wide range of navigation tasks primarily through visual sensory input and the recent success of vision-only navigation methods (Shah et al., 2023a; Zeng et al., ), we formulate the generalist navigation task as processing egocentric videos (captured by one or more cameras mounted on the robot) alongside language instructions, and predicting subsequent trajectories to fulfill those instructions. This formulation is compatible with most existing navigation task settings (Contributors, 2023; Wang et al., 2024a).
To align generalizable embodiments across diverse camera configurations, we introduce temporal-viewpoint indicator tokens (TVI tokens) to identify both the viewpoint of camera setups and the temporal information of the navigation horizon. By dynamically adjusting these TVI tokens, our method enables co-tuning across different camera setups and supports joint training with both image-QA and video-QA samples (Shen et al., 2024; Li et al., 2023). Furthermore, to address the constraints of practical deployment such as hardware memory cost and inference speed, we propose a token Budget-Aware Temporal Sampling (BATS) strategy, which dynamically samples navigation history tokens based on a forgetting curve constrained by a token budget. This token sampling approach balances performance and inference speed, enhancing the practicality of our method for real-world deployment.
@@
Our experiments demonstrate that NavFoM achieves substantial advancements in generalist navigation. Without task-specific fine-tuning, NavFoM attains state-of-the-art or competitive performance across diverse public benchmarks for a variety of embodiments. On VLN-CE RxR (Ku et al., 2020), NavFoM improves performance in multi-camera settings (from $56.3\%$ to $64.4\%$ SR) and in single-camera settings (from $51.8\%$ to $57.4\%$ SR) compared to prior baselines. On HM3D-OVON (Yokoyama et al., 2024b), our method achieves $45.2\%$ SR in a zero-shot setting, outperforming the previous fine-tuned SOTA method ($43.6\%$ SR). Similarly strong results are observed across various benchmarks in object searching, tracking, and autonomous driving. We further validate NavFoM through real-world experiments on multiple robotic platforms, including humanoid robots, quadrupeds, drones, and wheeled robots. These results underscore its strong generalizability and highlight promising progress toward generalist navigation.
@@ 2 Related Works
Large models for Navigation. Integrating Large Models (LLMs and VLMs) into robotic navigation has shifted the field from traditional learning-based methods toward leveraging pre-trained knowledge for open-world understanding and strong generalization capability. One straightforward approach (Zhou et al., 2023; Shah et al., 2022; Qiao et al., 2023) is to use off-the-shelf LLMs in a zero-shot manner. These works emphasize interpretability via chain-of-thought mechanisms (Pan et al., 2023; Long et al., 2023; Lin et al., 2025) and structured reasoning frameworks (Chen et al., 2024b; Qiao et al., 2025). However, abstracting dense visual information into text results in sparse environmental observations and is limited to static environments. Another pathway (Cheng et al., 2025; Zhang et al., 2024a; 2025a; Wei et al., 2025; Wang et al., 2025c) involves end-to-end fine-tuning of video-based or image-based (Zhou et al., 2025; Zheng et al., 2024; Zhang et al., 2025c) vision-language models with navigation data to enable VLMs to master navigation capabilities. However, existing methods mostly focus on isomorphic embodiments, overlooking the potential training synergy between different embodiments and tasks. In this work, we make an initial effort to extend the navigation policy to a broader domain of cross-embodiment and cross-task navigation.
@@
Cross-embodiments Navigation. The development of navigation models that generalize across diverse embodiments—varying in shape, size, and sensor configurations—remains a significant challenge in embodied AI. Recent efforts (Shah et al., 2023a; b; Yang et al., 2024b; Wang et al., 2020; Eftekhar et al., 2024; Hirose et al., 2023; Putta et al., 2024; Curtis et al., 2024; Wang et al., 2025a; Zhang et al., 2025b) have demonstrated the potential of transformer-based policies trained on large-scale, cross-embodiment datasets to achieve robust performance across various robotic platforms without the need for manual alignment of observation and action spaces. However, these models often process multimodal inputs without incorporating explicit spatial and temporal cues, which can lead to ambiguities due to the differing geometric interpretations of data from various embodiments. This approach may also result in data inefficiencies and limited generalization to out-of-distribution embodiments. (Eftekhar et al. (2024); Wang et al. (2025b)) In contrast, NavFoM introduces temporal-spatial indicator tokens that encode the observation configurations, enabling the model to better interpret multimodal inputs across different embodiments.
@@
Cross-task Navigation. Recent advances in embodied AI (O’Neill et al., 2024; Team et al., 2024; Kim et al., ; Bjorck et al., 2025; Black et al., 2024; Intelligence et al., 2025; Bu et al., 2025b; a; Qu et al., 2025) demonstrate that generalist models built upon foundation models can effectively transfer knowledge across diverse tasks. In the domain of navigation, prior studies (Zhou et al., 2024; Wang et al., 2022; Long et al., 2024; Song et al., 2025; Zhang et al., 2025a; Gao et al., 2025; Yin et al., 2025; Ruan et al., 2025) have shown that integrating data from different categories of navigation tasks can lead to stronger performance across various navigation scenarios. An early work, VIENNA (Wang et al., 2022), leveraged reinforcement learning to train an agent in simulators. More recently, Uni-Navid (Zhang et al., 2025a) developed a generalist model based on video-based vision-language models (VLMs) that learns generalized navigation skills through cross-task learning across four types of tasks: vision-and-language navigation, object-goal navigation, embodied question answering, and human following. However, these approaches are limited to constrained settings (e.g., indoor controllable environments), whereas our work extends to broader scenarios (e.g., including autonomous driving and UAV navigation) and unifies all tasks under a general formulation. In this framework, the model takes RGB videos and natural language instructions as input and outputs an executable trajectory.
@@ 3.2 Data
Vision-and-Language Navigation (3.37 M) requires an agent to interpret natural language instructions and egocentric visual observations, align the instructions with visual inputs, and plan subsequent actions to reach described landmarks. Following a broad definition of VLN (Zheng et al., 2024; Wang et al., 2025b; Zhou et al., 2024), we consider both indoor environments (e.g., VLN-CE on R2R (Krantz et al., 2020) and RxR (Ku et al., 2020)) and outdoor environments (e.g., OpenUAV (Wang et al., 2024a)), deployed on robots and unmanned aerial vehicles (UAVs), respectively. For VLN-CE on R2R and RxR (2.94 M), we capture multi-view RGB videos, annotated instructions, and trajectory data while the robot follows the ground-truth path. The multi-view RGB setup consists of a fixed front-view camera and randomly sampled surrounding cameras (ranging from one to eight). Camera heights are randomized between 0.6 m and 1.5 m, and the horizontal fields of view (HFoV) vary between 75 and 120. For the OpenUAV dataset (429 K), we record camera streams from the front, left, right, and rear views for all sequences. Other randomization strategies remain consistent with those used in the VLN-CE tasks.
@@ 4.1 Experiment Setup
Metrics. To evaluate navigation performance, we follow the standard evaluation metrics (Anderson et al., 2018), including success rate (SR), oracle success rate (OS), success weighted by path length (SPL), normalized Dynamic Time Warping (nDTW) and navigation error from goal (NE). Espeecially for tracking task, we leverage tracking rate (TR) (Puig et al., 2023), which measure the ratio of success tracking steps over all steps. Note that the success criteria change among different navigation tasks, we therefore use the default success criteria of each benchmark. For autonomous driving evaluation, we report L2 distance and Collision Rate (CR) for open-loop planning (Caesar et al., 2020a). For closed-loop evaluation on NAVSIM, we use the PDM score (PDMS) (Dauner et al., 2024b), which is a weighted combination of several sub-metrics: no at-fault collisions (NC), drivable area compliance (DAC), time-to-collision (TTC), comfort (Comf.), and ego progress (EP).